### v3

@@ Abstract
This paper presents the first application of Gaussian Mixture Copula Models to the statistical modeling of driving scenarios for the safety validation of automated driving systems. Knowledge of the joint probability distribution of scenario parameters is essential for scenario-based safety assessment, where risk quantification depends on the likelihood of concrete parameter combinations. Gaussian Mixture Copula Models bring together the multimodal expressivity of Gaussian Mixture Models and the flexibility of copulas, enabling separate modeling of marginal distributions and dependencies. We benchmark Gaussian Mixture Copula Models against previously proposed approaches—Gaussian Mixture Models and Gaussian Copula Models—using real-world driving data drawn from two scenarios defined in United Nations Regulation No. 157. Our evaluation on approximately 18 million instances of these two scenarios demonstrates that Gaussian Mixture Copula Models consistently surpass Gaussian Copula Models and perform better than, or at least comparably to, Gaussian Mixture Models, as measured by both log-likelihood and Sinkhorn distance. These results are promising for the adoption of Gaussian Mixture Copula Models as a statistical foundation for future scenario-based validation frameworks.
@@
Gaussian mixture copula models, scenario-based safety assessment, automated driving, joint probability estimation
@@ 1 INTRODUCTION
Central to this approach is the ability to determine the joint probability of specific parameter combinations, as this directly underpins the calculation of risk and ensures that simulated scenarios are relevant for real-world operation. ISO 21448 defines risk as the “combination of the probability of occurrence of harm and the severity of that harm” [1, subclause 3.23]. Without precise knowledge of these probabilities, the residual risk cannot be effectively quantified, nor can the most critical scenarios for validation be prioritized [11].
@@
  • Gaussian Mixture Models (GMMs) approximate the joint density with a weighted sum of elliptical Gaussians [14] and have been successfully applied to driving scenarios, such as the highway cut-in risk analysis by Akagi et al. [15]. While GMMs can approximate almost any continuous distribution given a sufficiently large number of components [16], practical applications typically rely on finite mixtures with a constrained number of components. In this finite regime, their reliance on multivariate Gaussian components can be limiting in practice when capturing intricate dependencies between parameters: the Gaussian assumptions may oversimplify complex relationships and higher-order interactions, thereby limiting their ability to model non-linear dependencies effectively.
@@
In this paper, we explore Gaussian Mixture Copula Models (GMCMs) to overcome the limitations of the previously proposed models. GMCMs integrate the strengths of GMMs and copula theory by decoupling marginal and dependency modeling. While GMCMs have been proposed and analyzed in general machine learning contexts [19, 20, 21, 22, 23], they have not yet been applied to the domain of automated driving.
@@ A GMM describes the joint density of random variables as a weighted sum of $K$ normal distributions,
where $\alpha^{k}\geq 0$ denotes the weight of the $k$-th component, such that $\sum_{k=1}^{K}\alpha^{k}=1$. $\phi(x_{1},x_{2},\dots x_{d};\bm{\mu}^{k},\Sigma^{k})$ represents a multivariate normal distribution with mean vector $\bm{\mu}^{k}$ and covariance matrix $\Sigma^{k}$. The parameters $\Theta=\{\alpha^{k},\bm{\mu}^{k},\Sigma^{k}\}_{k=1}^{K}$ of a GMM can be efficiently learned using the expectation-maximization (EM) algorithm [24]. The two-step approach of the EM algorithm guarantees convergence to a local maximum of the likelihood function. As stated before, a limitation of GMMs is the assumption that each of the $K$ components follows a multivariate normal distribution. For finite mixtures, this assumption can become restrictive in practice when the data exhibit non-Gaussian characteristics, such as skewness, heavy-tailed distributions, or intricate dependencies between variables, as is the case for our scenario parameters.
@@ 4 EXPERIMENT
In this section, we present our experimental study of GMCMs for joint probability estimation of scenario parameters. Our experiments, conducted on two real-world datasets from certification-relevant scenarios, show that GMCMs consistently outperform GCMs. With comparable computational resources, GMCMs achieve higher accuracy than GMMs for one scenario and comparable performance for the other. These results confirm previously reported advantages of GMCMs [19, 23] and extend them to the domain of scenario parameter modeling.
@@ 4.1 REAL-WORLD SCENARIO DATASETS
    Deceleration: “the ‘other vehicle’ suddenly decelerates in front of the ‘ego vehicle”’[3, Annex 4, Appendix 3, para. 2.2(c)]
@@ 4.2 RESULTS
Figure 4: Bivariate marginal distribution of the original data and samples from GMCM, GMM, and GCM, illustrated via density heatmaps (colour scale: turbo). Visual inspection indicates that the GMCM fits the data more faithfully than the GMM and the GCM.
@@
Table 3: Log-likelihood, Sinkhorn distance, and median training wall-clock time achieved by GMCM, GCM, and GMM on our Follow Lead Vehicle Emergency Brake and Deceleration datasets. The best-performing model for each metric and dataset is shown in bold.
@@ 5 CONCLUSION
This paper presents the first investigation of GMCMs for estimating joint distributions of scenario parameters in the context of scenario-based validation for automated driving systems. We compared GMCMs with previously proposed models for this application—GMMs and GCMs. This comparison confirmed that the advantages reported in earlier studies by Tewari et al. [19, 23] generalize effectively to the domain of scenario parameter modeling for automated driving. GMCMs combine the ability of GMMs to represent multimodal distributions with the flexibility of copulas to model arbitrary marginal distributions independently of their joint dependencies.
@@
GMCMs demonstrated superior performance on one dataset and, on the other, performed comparably to GMMs (with a slight GMM edge) while exceeding the GCM, even without hyperparameter fine-tuning. This suggests that GMCMs offer a promising alternative to established approaches for probabilistic modeling of scenario parameters. Future work should explore the extent to which performance can be improved by tuning estimation parameters, such as selecting the optimal number of mixture components via the BIC and adjusting convergence criteria [44]. Moreover, GMCMs should be applied to other scenario types defined in UN R157, such as Cut-in and Cut-out. Their performance in higher-dimensional settings should also be evaluated, for example by incorporating parameters related to roadway and environmental conditions. Finally, we restricted our analysis to continuous parameters. How discrete scenario parameters could be integrated into the GMCM framework remains an open question and is left for future work.

### v4

@@ Abstract
This paper presents the first application of Gaussian Mixture Copula Models to the statistical modeling of driving scenarios for the safety validation of automated driving systems. Knowledge of the joint probability distribution of scenario parameters is essential for scenario-based safety assessment, where risk quantification depends on the likelihood of concrete parameter combinations. Gaussian Mixture Copula Models bring together the multimodal expressivity of Gaussian Mixture Models and the flexibility of copulas, enabling separate modeling of marginal distributions and dependence. We benchmark Gaussian Mixture Copula Models against previously proposed approaches—Gaussian Mixture Models and Gaussian Copula Models—using real-world driving data drawn from two scenarios defined in United Nations Regulation No. 157. Our evaluation on approximately 18 million instances of these two scenarios demonstrates that Gaussian Mixture Copula Models consistently surpass Gaussian Copula Models and perform competitively with Gaussian Mixture Models, as measured by both log-likelihood and Sinkhorn distance, with relative performance depending on the scenario. The results are promising for the adoption of Gaussian Mixture Copula Models as a statistical foundation for future scenario-based validation frameworks.
@@
Automated driving, Gaussian Mixture Copula Models (GMCMs), joint probability estimation, scenario-based safety assessment.
@@ 1 INTRODUCTION
Central to this approach is the ability to determine the joint probability of specific parameter combinations, as this directly underpins the calculation of risk and ensures that simulated scenarios are relevant for real-world operation. ISO 21448 defines risk as the “combination of the probability of occurrence of harm and the severity of that harm” [32, subclause 3.23]. Without precise knowledge of these probabilities, the residual risk cannot be effectively quantified, nor can the most critical scenarios for validation be prioritized [24].
@@
  • Gaussian Mixture Models (GMMs) approximate the joint density with a weighted sum of multivariate Gaussian components [5] and have been successfully applied to driving scenarios, such as the highway cut-in risk analysis by Akagi et al. [2]. GMMs are well known to be universal approximators of continuous distributions given a sufficiently large number of components [6]. In practical applications, however, the number of components is typically finite, which reflects a trade-off between flexibility and robustness. In this finite-mixture regime, the elliptical structure of individual components may lead to oversmoothing of localized or strongly non-Gaussian features, particularly when complex marginal effects dominate.
@@
In this paper, we explore Gaussian Mixture Copula Models (GMCMs) as an extension of mixture-based modeling that retains the dependence structure of GMMs while increasing flexibility through separate, potentially non-parametric, marginal modeling. Specifically, GMCMs use a Gaussian-mixture-based copula construction that decouples marginal and dependence modeling. While GMCMs have been proposed and analyzed in general machine learning contexts [40, 3, 4, 16, 41], they have not yet been applied to the domain of automated driving.
@@ A GMM describes the joint density of random variables as a weighted sum of $K$ normal distributions,
where $\alpha^{k}\geq 0$ denotes the weight of the $k$-th component, such that $\sum_{k=1}^{K}\alpha^{k}=1$. $\phi(x_{1},x_{2},\dots x_{d};\bm{\mu}^{k},\Sigma^{k})$ represents a multivariate normal distribution with mean vector $\bm{\mu}^{k}$ and covariance matrix $\Sigma^{k}$. The parameters $\Theta=\{\alpha^{k},\bm{\mu}^{k},\Sigma^{k}\}_{k=1}^{K}$ of a GMM can be efficiently learned using the expectation-maximization (EM) algorithm [11]. The two-step approach of the EM algorithm guarantees convergence to a local maximum of the likelihood function.
A defining property of GMMs is that each mixture component follows a multivariate normal distribution. While mixtures of such components can approximate highly complex joint distributions in principle, finite mixtures may smooth sharp marginal features or heavy-tailed behavior unless a sufficiently large number of components is used, consistent with characteristics observed to some extent in our scenario parameters. In many finite-sample settings, this smoothing can also be beneficial, as it mitigates overfitting and improves generalization.
@@ 4 EXPERIMENT
In this section, we present our experimental study of GMCMs for joint probability estimation of scenario parameters. Our experiments, conducted on two real-world datasets from certification-relevant scenarios, show that GMCMs consistently outperform GCMs. With comparable computational resources, GMCMs achieve higher accuracy than GMMs for one scenario and comparable performance for the other, indicating that both approaches are competitive and that performance depends on the specific data characteristics. These findings are consistent with previously reported results on GMCMs [40, 41] and extend them to the domain of scenario parameter modeling.
@@ 4.1 REAL-WORLD SCENARIO DATASETS
    Deceleration: “the ‘other vehicle’ suddenly decelerates in front of the ‘ego vehicle”’[43, Annex 4, Appendix 3, para. 2.2(c)]
@@ 4.2 RESULTS
Figure 4: Bivariate marginal distribution of the original data and samples from GMCM, GMM, and GCM, illustrated via density heatmaps (colour scale: turbo).
@@
Table 3: Log-likelihood, Sinkhorn distance, and median training wall-clock time achieved by GMCM, GCM, and GMM on our Follow Lead Vehicle Emergency Brake and Deceleration datasets. Labels of the form GMM-K and GMCM-K indicate the number of mixture components K. The best-performing model for each metric and dataset is shown in bold.
@@ 5 CONCLUSION
This paper presents the first investigation of GMCMs for estimating joint distributions of scenario parameters in the context of scenario-based validation for automated driving systems. We compared GMCMs with previously proposed models for this application—GMMs and GCMs. This comparison confirmed that the advantages reported in earlier studies by Tewari et al. [40, 41] generalize effectively to the domain of scenario parameter modeling for automated driving. GMCMs combine the ability of GMMs to represent multimodal distributions with the flexibility of copulas to model arbitrary marginal distributions independently of their dependence structure.
@@
GMCMs demonstrated improved performance over GMMs in one dataset and comparable performance (with a slight GMM edge) in the other, while consistently outperforming GCMs, even without hyperparameter fine-tuning. Importantly, GMCMs retain the dependence structure of GMMs. Their benefits arise from decoupling marginal and dependence modeling, which can be advantageous when accurate marginal modeling is particularly important, for example in the presence of heavy tails or sharp, localized marginal features. This suggests that GMCMs offer a complementary modeling alternative that emphasizes flexible marginal estimation, leading to a different bias–variance trade-off compared to standard GMMs.
Future work should explore the extent to which performance can be improved by tuning estimation parameters, such as selecting the optimal number of mixture components via the BIC and adjusting convergence criteria [33]. Moreover, GMCMs should be applied to other scenario types defined in UN R157, such as Cut-in and Cut-out. Their performance in higher-dimensional settings should also be evaluated, for example by incorporating parameters related to roadway and environmental conditions. Finally, we restricted our analysis to continuous parameters. How discrete scenario parameters could be integrated into the GMCM framework remains an open question and is left for future work.